5/x + 7/6 = 4/(x-4), solve for x.
I am confused as to how I come up with the LCD for these to be able to solve for x

Answer by stanbon(57250)

5/x + 7/6 = 4/(x-4)
--------
Since you have an equation you don't want an lcd, you want a least common
multiple.
lcm = 6x(x-4)
--------
Multiply thru by the lcm to get:
5*6(x-4) + 7x(x-4) = 4*6x
30x-120 + 7x^2-28 = 24x
7x^2 + 6x - 148 = 0
Solve with the quadratic formula
======================================
Cheers,
Stan H.